OpenStudy (anonymous):

Choose the system of equations which matches the following graph. x + 2y = −10 3x − y = 12 x − 2y = −10 3x − y = 12 x + 2y = −10 3x + y = 12 x − 2y = −10 3x + y = 12

OpenStudy (anonymous):

@reemii

OpenStudy (reemii):

if you look at the propositions your head will hurt.. just find the two equations. You know - how to find the slope of both lines - you can find a point of the first and the second line.. then you can use a formula to write the equations.

OpenStudy (anonymous):

How would I get started on doing this for these equations?

OpenStudy (reemii):



OpenStudy (anonymous):

Yes, (-4, -3)

OpenStudy (anonymous):

Right? @reemii

OpenStudy (reemii):

find another point, and you'll be able to determine the slope.

OpenStudy (anonymous):

(4, -7)?

OpenStudy (reemii):

yes. then the slope is: ... other question: do you know the formula for the equation of a line? (i think you should fine it in your notes somewhere)

OpenStudy (anonymous):

I could go back and look through a semester of notes, but could you just show me what the equation is so I can find it?

OpenStudy (reemii):

\(y=ax+b\). if you have the slope (here it is -1/2) then you have a. a is the slope. b is the y-value on the y-axis. you can see that your line crosses the y-axis at y=-5. therefore the first equation is y=-1/2x-5.

OpenStudy (anonymous):

That isn't any of the possible first equations in the possible answers though

OpenStudy (reemii):

multiply everything by 2. you get 2y = -x-10 which is x+2y=-10.

OpenStudy (reemii):

so it's either the 1st or the 3rd option.

OpenStudy (anonymous):

Then the second equation would be, 3x + y = 12?

OpenStudy (reemii):

why?

OpenStudy (anonymous):

To be honest with you, that was just a guess

OpenStudy (rajee_sam):

There is a much simpler way. Would you like to know?

OpenStudy (rajee_sam):

@kyleguitarplayer

OpenStudy (anonymous):

Yes please!

OpenStudy (rajee_sam):

Ok. Each system of equations has two equations right? Each equation represents a line. That is why you have two lines in the graph. Each line belongs to one equation in the system

OpenStudy (rajee_sam):

When you solve these equations you will get a value for x and a value for Y right?

OpenStudy (rajee_sam):

are you there? pl. nod if you understand what I wrote

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

Yes I understand!

OpenStudy (rajee_sam):

sorry lost connection

OpenStudy (rajee_sam):

Ok when you see the two lines they intersect at a point. Can you tell me what point that is?

OpenStudy (rajee_sam):

@kyleguitarplayer

OpenStudy (anonymous):

The solution? The intersection point?

OpenStudy (rajee_sam):

Are you there @kyleguitarplayer ?

OpenStudy (rajee_sam):

yes the intersection point is the solution. So plug in this point in the system of equation and see which pair it is solution to.

OpenStudy (rajee_sam):

that is all you need to do

OpenStudy (anonymous):

So its option A then???

OpenStudy (rajee_sam):

yes

OpenStudy (anonymous):

Thank you so much for helping me understand!!